A gear train of a 28 tooth pinion and a 60 tooth spur gear is designed to operate with the following:

a 25 degree pressure angle, full-depth involute teeth, and a diametral pitch of 16 teeth/in. The standard center distance is increased by 0.125 inches. At what pressure angle do the gears operate now?
Answer #1

module = 1/Diametrical pitch = 1/16 = 0.0625 inch

now initial centre disatnce = m*(Tp+Tg)/2 where Tp and Tg are tooth on pinion and gear respectively

0.0625*(28+60)/2 = 2.75 inch

new centre distance =2.75+0.125 = 2.825 inch

But base circle radius remains same

so 2.75 *cos 25 = 2.825 * cos

where is new pressure angle

so cos = 0.8822

= 28.085 degrees